Institut for Matematik og Datalogi Odense Universitet February 8, 1996

JFB

DM11 – F96 – Lecture 2

Lecture, February 7

We began with an introduction to the course. Then, we will covered most of sections 1.1 through 1.7 in the textbook.

Lecture, February 14

We will finish discussing cardinality, from section 1.7. Then, we will begin on chapter 2. In this chapter, we will cover sections 2.3-2.5.

Lecture, February 15

We will continue on chapter 2 and possibly begin chapter 3.

Problems to be discussed in week 8

The following could be a property of a set S of integers. Write down the formal negation of this property: ∀x ∈ S, x ≥ k : ∃y ∈ S : x^{3} = y. 10.e, and 26 from section 1.5. (Note the definition before problem 22.)

10, 22, 23, 26, 37, and 39 from section 1.6. (Note the definitions before problems 24 and 39.) 12, 13, 14, 20, 21, 24, 25, 27, 28, and 30 from section 1.7. (Read problem 11 before starting the problems from this section.) For problem 28, you must assume that not all of a, b, and c are equal to zero. Why?